OpenStudy (anonymous):

What is negative 1 over 2 x + 2 = −x + 7 ?

OpenStudy (anonymous):

\[(-\frac{ 2x+2 }{ 1 })-\frac{ 1 }{ 2x+2 }=-x+7(-\frac{ 2x+2 }{ 1})\] Can you finish it from here?

OpenStudy (anonymous):

Yes i can thank you sir.